# Vision Foundation Models as Effective Visual Tokenizers for Autoregressive Generation

### v1

@@ 5 Conclusion
Acknowledgments
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We are deeply grateful to Lufan Ma for the contribution in polishing up this paper. . We would also appreciate Tong Yang for providing the DINO discriminator script. Besides, we are grateful to Lufan Ma for polishing up the details of the paper.
Author Contribution Statement
X.Q. proposed the initial concept of region-adaptive quantization. Based on this, A.Z. built the VFMTok, carried out the overall experiments, and led the writing of the initial draft. X.W., X.Q., and C.M. were deeply involved in the project discussions and manuscript writing. X(iangyu).Z., X(uanyang).Z., G.Y., and T.W., provided sufficient computational resource. X(uanyang).Z. participated in discussion, suggested ablation studies along with T.W., and discussed the writing of the draft. All authors contributed critical feedback, shaping the research, analysis, and the final manuscript.
@@ Acknowledgments
This section first illustrates the implementation of VFMTok and AR generation models, covering their learning rates, optimization approaches, and training demands. Subsequently, we present more ablation studies on VFMTok, investigating the effects of different architectural designs on image reconstruction and generation. Besides, all VFMToks in this study utilize a frozen, pre-trained DINOv2-L [9] as their encoder. Next, we show AR image generation with the other VFMs. Finally, we demonstrate additional visualization samples of class-to-image generation using VFMTok.
@@ Appendix B Supplemental Ablation Study
In this subsection, we conduct more ablation studies on the design of our approach, including the AR image generation with resolution of $256\times 256$, effect of single-level v.s. multi-level features, effect of shared ViT v.s. unshared ViT, impact of different components, and the number of tokens to represent an image.
@@ B.1 AR Generation with Resolution of 256×256
In the main paper, given that the input resolution of the vision foundation model is $336\times 336$, we adjust the resolution of reconstructed and generated images to $336\times 336$ by default, thus avoiding changing the number of tokens for image representation. Following the common practices [43, 44], we also train the image tokenizer and the AR generation model with the resolution of $256\times 256$, respectively. We first initialize and train VFMTok tokenizer for 50 epochs, then integrate it with AR generative models. Considering computational costs, models with fewer parameters like VFMTok-B and VFMTok-L are trained for 300 epochs, while larger AR models for 50 epochs. This setting aligns with LlamaGen [43] for $256\times 256$ image generation. As shown in Tab. 5, VFMTok not only achieves a decent reconstruction performance but also improves the generation quality compared to its counterparts [43] by a large margin. It is worth noting that VFMTok also accelerates the convergence speed during AR model training and significantly improves synthesis quality.
@@
In this paragraph, we investigate the separate impacts of single-level or multi-level features on image reconstruction and generation. Specifically, we replace the multi-level features with the single-level feature from the last layer of a VFM.
Table 6: Single-level v.s. multi-level features.
Type Image recon. $\mathcal{Q}_{C}$ AR gen. L.P.
@@
single-level 256 1.23 214.8 100.0% 3.64 277.7 41.3
multi-level 256 0.89 215.4 100.9% 3.09 274.2 69.4
Setup. We initialize and train each tokenizer for 50 epochs. Once the tokenizer is optimized, it is integrated with an AR generation model, VFMTok-L, for a total of 100 training epochs. Besides, the image reconstruction and generation performance are reported. Additionally, linear probing is also conducted on the [CLS]monospace-$[CLS]+$] token to estimate the semantic representation capability of VFMTok.
Observation. As depicted in Tab. 6, compared to VFMTok-L trained with single-level feature, the one with multi-level features presents a significantly superior performance in both image reconstruction, generation, and semantic representation capability.
@@ B.4 Effect of Different Components
Setup. To answer the above questions, we conducted 3 different experiments: 1) ${M}_{\text{I}}$ and $\textbf{M}_{f}$ share the same mask token $\text{M}_{\text{share}}$; 2) Replacing deformable attention with cross-attention in deformable transformer. To address the memory, computational demands, and fairness considerations of cross-attention, we first concatenate multi-level features from a VFM along the channel dimension, then apply a single MLP for dimensionality reduction before these features are fed to the cross-attention transformer for interaction with queries. Besides, each query interacts with VFM features within a $16\times 16$ window to simulate region-adaptive behavior; 3) Removing the register tokens from VFMTok. Additionally, linear probing is carried out on the [CLS]monospace-$[CLS]+$] token to estimate the semantic representation capability of VFMTok.
@@ B.6 Effect of Different VFMs
Here, we explore employing different frozen VFMs [9, 57, 46] as the encoder of VFMTok. As shown in Tab. 10, VFMTok implemented with SigLIP [57] also consistently yields decent image reconstruction and generation quality based on AR models of different scales. Additionally, we also observe a potential correlation between the AR synthesis quality and different VFMs. VFM with stronger representation capabilities will achieve higher generation performance (last 3 rows of Tab. 10).

### v2

@@ 5 Conclusion
6 Acknowledgments
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We are deeply grateful to Lufan Ma for the contribution in polishing up this paper. We would also appreciate Tong Yang for providing the DINO discriminator script.
7 Author Contribution Statement
X.Q. proposed the initial concept of region-adaptive quantization. Based on this, A.Z. built the VFMTok, conducted the overall experiments, and led the writing of the initial draft. X.W., X.Q., and C.M. were deeply involved in the project progress and manuscript writing. X(iangyu).Z., X(uanyang).Z., G.Y., and T.W., provided sufficient computational resources. X(uanyang).Z. joint discussion where suggested ablation studies along with T.W., and discussed the writing of the draft. All authors contributed critical feedback, shaping the research, analysis, and the final manuscript.
@@ 7 Author Contribution Statement
This section first illustrates the implementation of VFMTok and AR generation models, covering their learning rates, optimization approaches, and training demands. Subsequently, we present more ablation studies on VFMTok, investigating the effects of different architectural designs on image reconstruction and generation. Besides, all VFMToks in this study utilize a frozen, pre-trained DINOv2-L [9] as their encoder. Next, we show AR image generation with the other VFMs. Subsequently, we answer the question of what makes VFMs a good visual tokenizer. Finally, we demonstrate additional visualization samples of class-to-image generation using VFMTok.
@@ Appendix B Supplemental Ablation Study
In this subsection, we conduct more ablation studies on the design of our approach, including the AR image generation with resolution of $256\times 256$, the effect of single-level v.s. multi-level features, the effect of shared ViT v.s. unshared ViT, the impact of different components, and the number of tokens to represent an image. Additionally, the effect of different VFMs on image reconstruction and generation is also presented.
@@ B.1 AR Generation with Resolution of 256×256
In the main paper, given that the input resolution of the vision foundation model is $336\times 336$, we adjust the resolution of reconstructed and generated images to $336\times 336$ by default, thus avoiding changing the number of tokens for image representation. Following the common practices [43, 44], we also train the image tokenizer and the AR generation model with the resolution of $256\times 256$, respectively. We first initialize and train VFMTok tokenizer for 50 epochs, then integrate it with AR generative models. Considering computational costs, models with fewer parameters like VFMTok-B and VFMTok-L are trained for 300 epochs, while larger AR models for 50 epochs. This setting aligns with LlamaGen [43] for $256\times 256$ image generation. Furthermore, to ensure a fair comparison with the advanced autoregressive generation framework, RAR [55], we also incorporated VFMTok into RAR [55], maintaining the same setup during the training phase. As shown in Tab.˜5, VFMTok not only achieves a decent reconstruction performance but also improves the generation quality compared to its counterparts [43, 55] by a large margin. It is worth noting that VFMTok also accelerates the convergence speed during AR model training and significantly improves synthesis quality.
@@
In this paragraph, we investigate the separate impacts of each single-level or multi-level features on image reconstruction and generation. Specifically, we begin by conducting an ablation study to evaluate the impact of each single-layer feature on image reconstruction and generation. Subsequently, we cumulatively add each feature layer to observe the combined effect.
Table 6: Performance of each single-level feature. $F_{i}$ represents the indexed feature level.
$F_{i}$ Image recon. AR gen.
@@
$F_{1}$ 256 1.04 186.4 100.0% 3.84 257.9
$F_{2}$ 0.95 200.6 3.79 272.7
$F_{3}$ 1.03 208.5 3.69 274.9
$F_{4}$ 1.23 214.8 3.64 277.7
Table 7: Performance of cumulatively added features. $F_{i}$ represents the indexed feature level.
$F_{i}$ Image recon. AR gen.
[…]
$F_{1}$ 256 1.04 186.4 100.0% 3.84 257.9
$F_{2}$ 0.94 205.0 3.69 274.4
$F_{3}$ 0.94 210.8 3.27 272.5
$F_{4}$ 0.89 215.4 3.09 274.2
Setup. We initialize and train each tokenizer for 50 epochs. Once the tokenizer is optimized, it is integrated with an AR generation model, VFMTok-L, for a total of 100 training epochs. During evaluation, the image reconstruction and generation performance are reported.
Observation. The results, as presented in the Table 7 and Table 7, show that while a single feature layer offers no significant advantage, the quality of both image reconstruction and generation markedly improves as more layers are incorporated.
@@ B.4 Effect of Different Components
Setup. To answer the above questions, we conducted 3 different experiments: 1) ${M}_{\text{I}}$ and $\textbf{M}_{f}$ share the same mask token $\text{M}_{\text{share}}$; 2) Replacing deformable attention with cross-attention in deformable transformer. To address the memory, computational demands, and fairness considerations of cross-attention, we first concatenate multi-level features from a VFM along the channel dimension, then apply a single MLP for dimensionality reduction before these features are fed to the cross-attention transformer for interaction with queries. Besides, each query interacts with VFM features within a $16\times 16$ window to simulate region-adaptive behavior; 3) Removing the register tokens from VFMTok. Additionally, linear probing is carried out on the $\verb+[CLS]$+ token to estimate the semantic representation capability of VFMTok.
@@ B.6 Effect of Different VFMs
Here, we explore employing different frozen VFMs [9, 58, 46] as the encoder of VFMTok, and incorporate them into different AR generative frameworks – LlamaGen [43] and RAR [55]. As shown in Tab. 11, VFMTok implemented with SigLIP [58] also consistently yields decent image reconstruction and generation quality based on AR models of different scales. Additionally, we also observe a potential correlation between the AR synthesis quality and different VFMs. VFM with stronger representation capabilities will achieve higher generation performance.
@@
RAR-L(SigLIP) 400 461M 256 1.46 6.18 312.9 0.78 0.64 2.26 5.78 204.3 0.79 0.62
RAR-XL(SigLIP) 955M 1.40 6.39 311.7 0.78 0.65 1.87 5.77 226.7 0.79 0.63
RAR-XXL(SigLIP) 1.5B 1.38 6.18 298.4 0.78 0.65 1.68 5.64 239.5 0.79 0.63
RAR-XL(SigLIP2) 400 461M 256 1.50 6.37 292.7 0.77 0.66 2.05 5.50 217.7 0.79 0.62
RAR-XL(SigLIP2) 955M 1.46 6.02 288.4 0.78 0.65 1.72 5.41 244.8 0.80 0.63
RAR-XXL(SigLIP2) 955M 1.43 6.07 291.0 0.79 0.65 1.65 5.44 266.5 0.81 0.63
B.7 What makes a VFM a good visual tokenizer?
To answer this problem, we first leverage existing vision foundation models, which were supervised with different learning objectives, e.g. masked image modeling in pixel or latent space (Pixel-MIM and Latent-MIM) and contrastive learning (C.L) – to construct VFMTok for image reconstruction and generation, and subsequently provide more insights on this question.
Setup. The encoder of vanilla VQGAN [43] is substituted with different vision foundation models (VFMs), which are supervised with distinct learning objectives. All of these tokenizers are trained on the ImageNet [10] training set for 50 epochs. Subsequently, these tokenizers are incorporated into an AR generative model – LlamaGen-L [43] and trained for 100 epochs. Both image reconstruction and generation quality are evaluated on the ImageNet validation set with FID and IS, respectively. Additionally, the top-1 accuracy of linear probing is also reported.
Observation As shown in the Table. 12, Pixel-MIM (MAE [15]) aligns best to reconstruction objectives, thus, VQGAN(MAE) achieves the optimal image reconstruction quality. However, it does not provide a feature space that is friendly for generation tasks. Its semantics is also inferior to the other VFMs. Contrastive learning-only models, VQGAN(CLIP) and VQGAN(SigLIP) achieve the best semantics (top-1 accuracy in linear probing), but modest image reconstruction and generation performance. Without a mask image modeling objective, they lack sufficient capability to model the local structures. Latent-MIM and contrastive learning co-optimized approaches, VQGAN(SigLIP2), and VQGAN(DINOv2) achieve the best overall performance.
Table 12: Image reconstruction and generation with VFM of a distinct learning objective.
[…]
VQGAN(CLIP [37]) 576 1.47 182.0 3.45 221.2 59.5
VQGAN(SigLIP [58]) 1.26 190.8 3.50 246.1 60.3
VQGAN(SigLIP2 [46]) 0.96 198.4 3.39 267.8 55.5
VQGAN(DINOv2 [9]) 0.99 206.3 3.34 268.6 56.4
VGQGAN(MAE [15]) 0.67 207.6 3.40 265.5 39.0
Discussion The training objectives of vision foundation models(VFMs) affect the image reconstruction and generation of VFMTok. CLIP [37] and SigLIP [58] apply contrastive learning to image-text pairs, whereas DINO [3] applies a contrastive learning-like clustering objective to images. MAE [15] performs mask image modeling on image patches in pixel space, and iBOT [62] predicts DINO-like cluster assignments of image patches in latent space. DINOv2 [9] basically is the combination of DINO [3] and iBOT [62], and SigLIP2 [46] is also essentially the integration of SigLIP [58] and iBOT [62] (termed TIPS loss in their paper), omitting other regularizing losses. We hereby provide a holistic comparison of the training objectives of different VFMs in Table 13.
Table 13: The category of learning objectives adopted in VFMs.
VFM P-MIM L-MIM C.L.
CLP [37]
DINO [3]
SigLIP [58]
MAE [15]
iBOT [62]
SigLIP2 [46]
DINOv2 [9]
Notably, iBOT [62], DINOv2 [9], and SigLIP2 [46] are performing codebook learning resemble that in VQGAN on image patches, despite codes being soft instead of one-hot, and their codebook vectors being high-dimensional (e.g., 256). Following DINO [3], they learn a set of 8,192 or 65,536 prototypes, and require two augmented versions of the same patch (or image) to have identical prototype assignments (soft codes). They also predict the soft codes of masked patches, thus performing mask image modeling(Latent-MIM) in latent space. This formulation is termed self-distillation in DINO [3], and then extended to patches with Laent-MIM by iBOT [62], which was subsequently inherited by DINOv2 [9] and SigLIP2 [46].
The connection between Latent-MIM and VQGAN makes it natural to convert these VFMs to tokenizers and expect decent image reconstruction and generation performance. Additionally, the global (image-level) contrastive learning objectives of DINOv2 [9] and SigLIP2 [46] also promise better high-level semantics (better performance on understanding tasks).
To summarize, the conclusion can be drawn as: masked image modeling(MIM) objectives primarily help reconstruction and generation. MIM in pixel space helps reconstruction more, but is less beneficial for generation than MIM in latent space. Contrastive learning(C.L.) objectives are less helpful for reconstruction and generation, but are important for understanding abilities (e.g., top-1 accuracy on ImageNet [10]). Best VFMs for visual tokenization are trained with both mask image modeling in latent space and contrastive objectives, namely DINOv2 [9] and SigLIP2 [9].